Calculate the molar mass for the precipitate that will be

formed, MgNH4PO4.6H2O.
Answer #1

The moler mass of Mg NiHAP04-6420 Da = 24.3050 + 14.0067 + 16x(1.00794) + 30.973762+ 10X(15.9994) = 245.41